how many moles of cadmium chloride can be produced from 45,1 grams of cadmium with excess sulfur?


1
Expert's answer
2022-04-14T12:17:03-0400

(1): How many moles of cadmium chloride can be produced from 45.1 grams of cadmium with excess chlorine gas?

Solution (1):

The molar mass of cadmium (Cd) is 112.411 g/mol

Therefore,

Moles of Cd = (45.1 g Cd) × (1 mol Cd / 112.411 g Cd) = 0.4012 mol Cd


Balanced chemical equation:

Cd + Cl2 → CdCl2

According to stoichiometry:

1 mol of Cd produces 1 mol of CdCl2

Thus, 0.4012 mol of Cd produce:

(0.4012 mol Cd) × (1 mol CdCl2 / 1 mol Cd) = 0.4012 mol CdCl2


The molar mass of cadmium chloride (CdCl2) is 183.32 g/mol

Therefore,

Mass of CdCl2 = (0.4012 mol CdCl2) × (183.32 g CdCl2 / 1 mol CdCl2) = 73.55 g CdCl2

Mass of CdCl2 = 73.55 g


Answer (1): 73.55 grams of cadmium chloride (CdCl2) are produced




(2): How many moles of cadmium sulfide can be produced from 45.1 grams of cadmium with excess sulfur?

Solution (2):

The molar mass of cadmium (Cd) is 112.411 g/mol

Therefore,

Moles of Cd = (45.1 g Cd) × (1 mol Cd / 112.411 g Cd) = 0.4012 mol Cd


Balanced chemical equation:

Cd + S → CdS

According to stoichiometry:

1 mol of Cd produces 1 mol of CdS

Thus, 0.4012 mol of Cd produce:

(0.4012 mol Cd) × (1 mol CdS / 1 mol Cd) = 0.4012 mol CdS


The molar mass of cadmium sulfide (CdS) is 144.47 g/mol

Therefore,

Mass of CdS = (0.4012 mol CdS) × (144.47 g CdS / 1 mol CdS) = 57.96 g CdS

Mass of CdS = 57.96 g


Answer (2): 57.96 grams of cadmium sulfide (CdS) are produced